The L and C State Building is 1, 250 m tall. What is the angle of elevation of 

the top from a point on the ground 1,760 m from the base of the building?


1
Expert's answer
2021-01-26T03:08:38-0500

Let our b be the base of the building (adjacent), a to be the height of the building(opposite), and lastly our c to be the elevation of the building ( hypotenuse).

We are asked for the angle of elevation of the top point, which is the angle B.

  • As we know SinA=a/b=Opposite/Hypotenuse
  • SinA =1250m/1760m
  • SinA = 0.7102
  • Sin-1 = 50.28o
  • 50.28o = Angle of A
  • 90.00o = Angle of C
  • As we know angles in a triangle add up to 180o
  • Angle A + Angle B + Angle C = 180o
  • 50.28o + 90.00o + X = 180o
  • X = 180 - (50.28o + 90.00o)
  • X = 39.72o.
  • The angle of elevation of the top from a point on the ground is 39.72o.